# SKD-TSTSAN: Three-Stream Temporal-Shift Attention Network Based on Self-Knowledge Distillation for Micro-Expression Recognition

### v1

@@ 3.2 Efficient Channel Attention Module
Therefore, we employ efficient channel attention (ECA) modules [14] in the L-stream, introducing an efficient channel attention mechanism. Specifically, compared to the original squeeze-and-excitation (SE) block [35], the ECA module avoids disrupting the direct relationship between channels and their weights and uses 1D convolution to achieve local cross-channel interaction. As illustrated in Fig. 2, the input feature map is $X\in{{\mathbb{R}}^{W\times H\times C}}$, where $W,H,C$ represents the width, height, and channel dimensions. The original ECA module employs global average pooling (GAP) to aggregate global spatial features into ${{F}_{avg}}\in{{\mathbb{R}}^{1\times 1\times C}}$. Additionally, we employ global maximum pooling (GMP) to provide the most significant information for each channel, obtaining ${{F}_{max}}\in{{\mathbb{R}}^{1\times 1\times C}}$. Then, by performing a 1D convolution with a
@@
kernel size of $k$ on ${F}_{avg}$ and ${F}_{max}$, we capture cross-channel interactions and obtain the channel weights $M\in{{\mathbb{R}}^{1\times 1\times C}}$, which are represented as follows:
@@ 3.3 Temporal Shift Module
3D CNNs can effectively mine spatiotemporal information but are computationally intensive. In order to extract spatiotemporal information while
keeping the complexity of 2D CNNs, we incorporate the temporal shift module (TSM) [15] into the T-stream.
The schematic of the TSM we use is shown in Fig. 3. Since the ME videos are offline videos, with access to future frames, the bi-directional TSM is employed. For a tensor with C channels and T frames, we use varying colors to represent features at different time stamps. By shifting part of the channels forward by one frame and another part backward by one frame along the temporal dimension, we achieve the mixing of information from past and future frames with the current frame. The excess channels are truncated and the blank frame parts are filled with zeros. The shift operation in TSM does not require any multiplication, and the multiply-accumulate operation of features from different frames’ channels is completed by the subsequent 2D convolution. Therefore, efficient temporal modeling is achieved at no extra cost.
As shown in Fig. 4(a), in the original TSCNN, the input to the T-stream is the concatenation of horizontal and vertical optical flows between the onset frame, apex frame, and offset frame pairs. It can be seen that 2D CNN without TSM have difficulty in capturing temporal relationships in forward propagation. To introduce TSM, we separate the inputs of the T-stream along the temporal dimension, i.e., the optical flow maps between the onset frame and the apex frame and between the apex frame and the offset
@@
frame, containing motion information from two different temporal domains. As illustrated in Fig. 4(b), TSMs are added before all convolutional layers in the intermediate layers of the T-stream to mix information from the two different temporal domains, enhancing the model’s temporal modeling capability. When the ratio of shifted channels in TSM is too big, it may affect the learning of spatial features; if too small, the temporal modeling capability is insufficient to handle complex temporal relationships. [15] Therefore, we shift 1/8 of the channels in the TSM.
@@ 3.4 Self-Knowledge Distillation
In order to further improve our lightweight three-stream network’s performance, we introduce an effective model compression strategy, self-knowledge distillation (SKD) [17]. In traditional knowledge distillation, a complex teacher model is first trained, and then a compact student model imitates the predictions of the teacher model [33]. However, it is difficult to design an effective over-parameterized model to act as the teacher model on a small-scale ME dataset [17]. Instead, in SKD, the model acts as its own teacher, using knowledge acquired from the training set to guide its own training. This approach requires less training time than the two-stage method and achieves better results.
We adopt the Be Your Own Teacher (BYOT) framework proposed by Zhang et al. [17], adding two auxiliary classifiers (ACs) after the first and third convolutional layers of the network. Each AC has a three-stream architecture similar to the main network, including convolutional layers and a fully connected layer, and contains the previously mentioned ECA modules and TSMs. The ACs are only used during training and can be removed during inference. During the training stage, the deepest part of the network acts as the teacher model, while the shallow parts with the corresponding classifiers serve as the student model. We employ three kinds of loss functions: focal loss (FL) [36], Kullback-Leibler (KL) divergence loss, and L2 loss. Then the computation methods for each loss function will be detailed.
@@ 4.1 Datasets
We evaluate our SKD-TSTSAN on four spontaneous ME datasets: the CASME II [7], SAMM [8], MMEW [9], and CAS(ME)3 [10] databases. The details of these datasets are introduced below.
@@ 4.2 Data Pre-Processing and Implementation Details
As the proposed SKD-TSTSAN requires knowledge transfer from macro-expression data, we first pre-train it on the CK+ dataset [16], which includes seven expressions: Happiness, Anger, Disgust, Contempt, Fear, Sadness, and Surprise. The CK+ dataset, being a macro-expression dataset, lacks annotations like the apex frames present in ME datasets. Therefore, to pre-train the network using this dataset, we designate the last frame with the most obvious muscle movement of each sample as the apex frame, and the first frame as both the onset and offset frames.
For each input sample, Retinaface [37] is first employed to detect the facial region and resize it to $128\times 128$. Then, we compute optical flow features using the TV-L1 method [38]. The entire facial region is scaled down to $48\times 48$ and fed into the S-stream of our network. The $128\times 128$ facial region is divided into a $4\times 4$ grid, resulting in 16 patches, each scaled to $48\times 48$. These patches are concatenated along the channel dimension and fed into the L-stream. Finally, the horizontal and vertical optical flow fields between the three key-frames are combined into two $48\times 48$ 2-channel images, which are fed into the T-stream of the network.
We implement our SKD-TSTSAN model on an NVIDIA-RTX-4090 GPU. The Adam optimizer is utilized with learning rate, betas and weight decay set to 1e-3, (0.9, 0.99) , 5e-4, respectively. Loss weights ${{\lambda}_{1}}=0.1$, ${{\lambda}_{2}}=1e-6$ and distillation temperature $T=3$ are chosen as default values. In the multi-class focal loss, the focusing parameter $\gamma$ is set to 2, and the weighting factor ${{\alpha}_{k}}$ is positively correlated with the inverse of the sample size for each category. In our motion magnification module, we set the amplification factor to 2. During the pre-training stage on the CK+ dataset, we configure the training iterations to 10k and the batch size to 16. When training on the CASME II, SAMM, and MMEW datasets with smaller data volumes, the batch size is set to 16, while for the CAS(ME)3 dataset, it is set to 32. Training iterations on the ME dataset are set to 20k and early stopping is used to alleviate overfitting.
@@ 4.4 Comparative Experiments Results
As shown in Table 2, we report the performance of SKD-TSTSAN on the CASME II dataset and compare it with other methods. Our SKD-TSTSAN achieves optimal performance on both 3-category and 5-category MER tasks. When performing the 3-category classification, SKD-TSTSAN achieves a superior 99.13% UF1 score and 98.67% UAR, exhibiting a 3.81% UF1 improvement and 3.51% UAR improvement compared to the second-ranked MER method, HTNet [23]. For the 5-category emotion classification, SKD-TSTSAN achieved the highest UF1 of 90.34% and UAR of 89.47%,
Table 1: Comparisons with other approaches on CASME II dataset.
Method # Classes UF1 (%) UAR (%)
LR-GACNN [39] 5 70.90 81.30
AMAN [26] 5 71.00 75.40
Graph-TCN [24] 5 72.46 73.98
MiMaNet [12] 5 75.90 79.90
SMA-STN [40] 5 79.46 82.59
TSCNN [11] 5 80.70 80.97
$\mu$-BERT [41] 5 85.53 83.48
SKD-TSTSAN (ours) 5 90.34 89.47
STSTNet [42] 3 83.82 86.86
OFF-ApexNet [43] 3 86.97 88.28
$\mu$-BERT [41] 3 90.34 89.14
HTNet [23] 3 95.32 95.16
SKD-TSTSAN (ours) 3 99.13 98.67
@@
OFF-ApexNet [43] 3 54.09 53.92
STSTNet [42] 3 65.88 68.10
EMR [44] 3 77.54 71.52
FeatRef [45] 3 73.72 71.55
HTNet [23] 3 81.31 81.24
@@
compared to $\mu$-BERT’s 85.53% UF1 and 83.48% UAR. Furthermore, compared to our baseline TSCNN, SKD-TSTSAN shows a 9.64% UF1 and 8.50% UAR improvement, proving the effectiveness of our submodules.
Table 2 presents the quantitative comparison results for the 3-category MER task on the SAMM dataset. Our SKD-TSTSAN achieves a UF1 of 94.29% and a UAR of 92.65%, showing an increase of 12.98% and 11.41% compared to the previous leading method HTNet [23]. All these methods are deep learning approaches that utilize optical flow features. The difference is that our SKD-TSTSAN uses the optical flow fields between the three key-frames as the input to the T-stream, while the other methods [42, 23, 43, 44, 45] only utilize the optical flow field between the onset frame and the apex frame. Therefore, SKD-TSTSAN can leverage motion information from different temporal regions for temporal modeling, fully considering the temporal information of ME and achieving the best recognition performance.
The performance of SKD-TSTSAN on the MMEW dataset is compared
@@
CapsuleNet [46] 6 44.13 45.31
MMNet [47] 6 57.05 56.95
STSTNet [42] 6 64.13 62.65
TSCNN [11] 6 70.43 68.76
@@
CapsuleNet [46] 4 64.02 63.15
STSTNet [42] 4 78.99 78.89
TSCNN [11] 4 86.46 86.09
MMNet [47] 4 86.35 87.45
@@
FR [45] 3 34.93 34.13
STSTNet [42] 3 37.95 37.92
RCN-A [48] 3 39.28 38.93
$\mu$-BERT [41] 3 56.04 61.25
@@
AlexNet [10] 4 29.15 29.10
AlexNet(+Depth) [10] 4 30.01 29.82
SFAMNet [49] 4 44.62 47.97
$\mu$-BERT [41] 4 47.18 49.13
@@
AlexNet [10] 7 17.59 18.01
AlexNet(+Depth) [10] 7 17.73 18.29
SFAMNet [49] 7 23.65 23.73
$\mu$-BERT [41] 7 32.64 32.54
@@
with different methods in Table 4. We conduct experiments on the MMEW dataset using CapsuleNet [46], MMNet [47], STSTNet [42], and TSCNN [11]. For MMNet’s performance on the 4-category MER task, the optimal performance reported in their original paper is provided. For six categories, CapsuleNet obtains the lowest UF1 and UAR since it only uses the apex frame as the network input, which makes it difficult to extract the motion and temporal information of MEs. In contrast, STSTNet and TSCNN use the optical flow as input and achieve better performance. Our SKD-TSTSAN achieves the best recognition performance, reaching 77.11% UF1 and 74.77% UAR and showing improvements of 6.68% and 6.01% over TSCNN. Similar
@@
improvements are observed in the 4-category emotion classification: a 3.95% over TSCNN for UF1 and a 2.99% increase over MMNet for UAR.
To further validate our proposed SKD-TSTSAN, we perform comparative experiments on the large-scale ME dataset CAS(ME)3 using 3, 4, and 7 emotion categories, as shown in Table 4. When tested with 3, 4, and 7 emotion categories, SKD-TSTSAN exhibits double-digit improvements over all other methods. In the case of 3 emotion categories, SKD-TSTSAN obtains a UF1 score of 86.48% and a UAR score of 83.44%, which indicates an improvement of 30.44% in UF1 and 22.19% in UAR compared to the second-ranked $\mu$-BERT [41]. For the 4 emotion categories, our SKD-TSTSAN maintains a considerable advantage over other methods with over 29.22% in UF1 and 23.44% in UAR. Using all 7 emotion categories is more demanding on the recognition performance of the MER method. SKD-TSTSAN significantly outperforms $\mu$-BERT in terms of UF1 (59.95% vs 32.64%) and UAR (57.31% vs 32.54%).
@@
datasets, SKD-TSTSAN achieves good recognition results for all categories. On the MMEW dataset, SKD-TSTSAN performs well on the 4-class MER task. On the 6-class MER task, SKD-TSTSAN has poor recognition performance for the “Fear” and “Sadness” samples. One of the main reasons is the small number of samples for the “Fear” and “Sadness” emotion categories. Besides, different emotion categories may correspond to the same AU. For example, both “Fear” and “Surprise” expressions appear in AU5 (“Upper Lid Raiser”), while “Fear” and “Disgust” expressions both involve AU4 (“Brow Lowerer”), as shown in Fig. 6. Therefore, samples labeled as “Fear” are easily misclassified as “Surprise” and “Disgust” categories, which have larger sample sizes. On the CAS(ME)3 dataset, the performance of SKD-TSTSAN is also affected by the imbalanced class distribution. For the 3 and 4 emotion categories, SKD-TSTSAN has poorer recognition performance on “Positive”, which has fewer samples. For the 7 emotion categories, SKD-TSTSAN performs better in the “Disgust”, “Others”, and “Surprise” categories. One of the reasons for the poor performance on the CAS(ME)3 dataset is that the onset frame annotations of some samples are the same as the apex frame annotations. This results in no motion information in the optical flow between them, leading to misclassification.

### v2

@@ 3.2 Efficient Channel Attention Module
Therefore, we employ efficient channel attention (ECA) modules (Wang et al.,, 2020) in the L-stream, introducing an efficient channel attention mechanism. Specifically, compared to the original squeeze-and-excitation (SE) block (Hu et al.,, 2018), the ECA module avoids disrupting the direct relationship between channels and their weights and uses 1D convolution to achieve local cross-channel interaction. As illustrated in Fig. 2, the input feature map is $X\in{{\mathbb{R}}^{W\times H\times C}}$, where $W,H,C$ represents the width, height, and channel dimensions. The original ECA module employs global average pooling (GAP) to aggregate global spatial features into ${{F}_{avg}}\in{{\mathbb{R}}^{1\times 1\times C}}$. Additionally, we employ global maximum pooling (GMP) to provide the most significant information for each channel, obtaining ${{F}_{max}}\in{{\mathbb{R}}^{1\times 1\times C}}$. Then, by performing a 1D convolution with a kernel size of $k$ on ${F}_{avg}$ and ${F}_{max}$, we capture cross-channel interactions and obtain the channel weights $M\in{{\mathbb{R}}^{1\times 1\times C}}$, which are represented as follows:
@@ 3.3 Temporal Shift Module
3D CNNs can effectively mine spatiotemporal information but are computationally intensive. In order to extract spatiotemporal information while keeping the complexity of 2D CNNs, we incorporate the temporal shift module (TSM) (Lin et al.,, 2019) into the T-stream.
The schematic of the TSM we use is shown in Fig. 3. Since the ME videos are offline videos, with access to future frames, the bi-directional TSM is employed. For a tensor with C channels and T frames, we use varying colors to represent features at different time stamps. By shifting part of the channels forward by one frame and another part backward by one frame along the temporal dimension, we achieve the mixing of information from past and
@@
future frames with the current frame. The excess channels are truncated and the blank frame parts are filled with zeros. The shift operation in TSM does not require any multiplication, and the multiply-accumulate operation of features from different frames’ channels is completed by the subsequent 2D convolution. Therefore, efficient temporal modeling is achieved at no extra cost.
As shown in Fig. 4(a), in the original TSCNN, the input to the T-stream is the concatenation of horizontal and vertical optical flows between the onset frame, apex frame, and offset frame pairs. It can be seen that 2D CNN without TSM have difficulty in capturing temporal relationships in forward propagation. To introduce TSM, we separate the inputs of the T-stream along the temporal dimension, i.e., the optical flow maps between the onset frame and the apex frame and between the apex frame and the offset frame, containing motion information from two different temporal domains. As illustrated in Fig. 4(b), TSMs are added before all convolutional layers in the intermediate layers of the T-stream to mix information from the two different temporal domains, enhancing the model’s temporal modeling capability. When the ratio of shifted channels in TSM is too big, it may affect the learning of spatial features; if too small, the temporal modeling capability is insufficient to handle complex temporal relationships (Lin et al.,, 2019). Therefore, we shift 1/8 of the channels in the TSM.
@@ 3.4 Self-Knowledge Distillation
In order to further improve our lightweight three-stream network’s performance, we introduce an effective model compression strategy, self-knowledge distillation (SKD) (Zhang et al.,, 2019). In traditional knowledge distillation, a complex teacher model is first trained, and then a compact student model imitates the predictions of the teacher model (Hinton et al.,, 2015). However, it is difficult to design an effective over-parameterized model to act as the teacher model on a small-scale ME dataset (Zhang et al.,, 2019). Instead, in SKD, the model acts as its own teacher, using knowledge acquired from the training set to guide its own training. This approach requires less training time than the two-stage method and achieves better results.
We adopt the Be Your Own Teacher (BYOT) framework proposed by Zhang et al., (2019), adding two auxiliary classifiers (ACs) after the first and third convolutional layers of the network. Each AC has a three-stream architecture similar to the main network, including convolutional layers and a fully connected layer, and contains the previously mentioned ECA modules and TSMs. The ACs are only used during training and can be removed during inference. During the training stage, the deepest part of the network acts as the teacher model, while the shallow parts with the corresponding classifiers serve as the student model. We employ three kinds of loss functions: focal loss (FL) (Lin et al.,, 2017), Kullback-Leibler (KL) divergence loss, and L2 loss. Then the computation methods for each loss function will be detailed.
@@ 4.1 Datasets
We evaluate our SKD-TSTSAN on four spontaneous ME datasets: the CASME II (Yan et al.,, 2014), SAMM (Davison et al.,, 2016), MMEW (Ben et al.,, 2021), and CAS(ME)3 (Li et al., 2022b, ) databases. The details of these datasets are introduced below.
@@ 4.2 Data Pre-Processing and Implementation Details
As the proposed SKD-TSTSAN requires knowledge transfer from macro-expression data, we first pre-train it on the CK+ dataset (Lucey et al.,, 2010), which includes seven expressions: Happiness, Anger, Disgust, Contempt, Fear, Sadness, and Surprise. The CK+ dataset, being a macro-expression dataset, lacks annotations like the apex frames present in ME datasets. Therefore, to pre-train the network using this dataset, we designate the last frame with the most obvious muscle movement of each sample as the apex frame, and the first frame as both the onset and offset frames.
For each input sample, Retinaface (Deng et al.,, 2019) is first employed to detect the facial region and resize it to $128\times 128$. Then, we compute optical flow features using the TV-L1 method (Zach et al.,, 2007). The entire facial region is scaled down to $48\times 48$ and fed into the S-stream of our network. The $128\times 128$ facial region is divided into a $4\times 4$ grid, resulting in 16 patches, each scaled to $48\times 48$. These patches are concatenated along the channel dimension and fed into the L-stream. Finally, the horizontal and vertical optical flow fields between the three key-frames are combined into two $48\times 48$ 2-channel images, which are fed into the T-stream of the network.
We implement our SKD-TSTSAN on an NVIDIA-RTX-4090 GPU. The Adam optimizer is utilized with learning rate, betas and weight decay set to 1e-3, (0.9, 0.99) , 5e-4, respectively. Loss weights ${{\lambda}_{1}}=0.1$, ${{\lambda}_{2}}=1e-6$ and distillation temperature $T=3$ are chosen as default values. In the multi-class focal loss, the focusing parameter $\gamma$ is set to 2, and the weighting factor ${{\alpha}_{k}}$ is positively correlated with the inverse of the sample size for each category. In our motion magnification module, we set the amplification factor to 2. During the pre-training stage on the CK+ dataset, we configure the training iterations to 10k and the batch size to 16. When training on the CASME II, SAMM, and MMEW datasets, the batch size is set to 16, while for the CAS(ME)3 dataset, it is set to 32. Training iterations on the ME dataset are set to 20k and early stopping is used to alleviate overfitting.
@@ 4.3 Evaluation Metrics
Table 1: Comparisons with other approaches on CASME II dataset.
Method # Classes UF1 (%) UAR (%)
LR-GACNN (Kumar and Bhanu,, 2021) 5 70.90 81.30
AMAN (Wei et al.,, 2022) 5 71.00 75.40
Graph-TCN (Lei et al.,, 2020) 5 72.46 73.98
MiMaNet (Xia and Wang,, 2021) 5 75.90 79.90
SMA-STN (Liu et al.,, 2020) 5 79.46 82.59
TSCNN (Song et al.,, 2019) 5 80.70 80.97
$\mu$-BERT (Nguyen et al.,, 2023) 5 85.53 83.48
SKD-TSTSAN (ours) 5 90.34 89.47
STSTNet (Liong et al.,, 2019) 3 83.82 86.86
OFF-ApexNet (Gan et al.,, 2019) 3 86.97 88.28
$\mu$-BERT (Nguyen et al.,, 2023) 3 90.34 89.14
HTNet (Wang et al.,, 2023) 3 95.32 95.16
SKD-TSTSAN (ours) 3 99.13 98.67
@@ 4.4 Comparative Experiments Results
As shown in Table 1, we report the performance of SKD-TSTSAN on the CASME II dataset and compare it with other methods. Our SKD-TSTSAN achieves optimal performance on both 3-category and 5-category MER tasks. When performing the 3-category classification, SKD-TSTSAN achieves a superior 99.13% UF1 score and 98.67% UAR, exhibiting a 3.81% UF1 improvement and 3.51% UAR improvement compared to the second-ranked MER method, HTNet (Wang et al.,, 2023). For the 5-category emotion classification, SKD-TSTSAN achieved the highest UF1 of 90.34% and UAR of 89.47%, compared to $\mu$-BERT’s 85.53% UF1 and 83.48% UAR. Furthermore, compared to our baseline TSCNN, SKD-TSTSAN shows a 9.64% UF1 and 8.50% UAR improvement, proving the effectiveness of our submodules.
Table 3 presents the quantitative comparison results for the 3-category MER task on the SAMM dataset. Our SKD-TSTSAN achieves a UF1 of
@@
OFF-ApexNet (Gan et al.,, 2019) 3 54.09 53.92
STSTNet (Liong et al.,, 2019) 3 65.88 68.10
EMR (Liu et al.,, 2019) 3 77.54 71.52
FeatRef (Zhou et al.,, 2022) 3 73.72 71.55
HTNet (Wang et al.,, 2023) 3 81.31 81.24
@@
CapsuleNet (Van Quang et al.,, 2019) 6 44.13 45.31
MMNet (Li et al., 2022a, ) 6 57.05 56.95
STSTNet (Liong et al.,, 2019) 6 64.13 62.65
TSCNN (Song et al.,, 2019) 6 70.43 68.76
@@
CapsuleNet (Van Quang et al.,, 2019) 4 64.02 63.15
STSTNet (Liong et al.,, 2019) 4 78.99 78.89
TSCNN (Song et al.,, 2019) 4 86.46 86.09
MMNet (Li et al., 2022a, ) 4 86.35 87.45
@@
94.29% and a UAR of 92.65%, showing an increase of 12.98% and 11.41% compared to the previous leading method HTNet (Wang et al.,, 2023). All these methods are deep learning approaches that utilize optical flow features. The difference is that our SKD-TSTSAN uses the optical flow fields between the three key-frames as the input to the T-stream, while the other methods (Liong et al.,, 2019; Wang et al.,, 2023; Gan et al.,, 2019; Liu et al.,, 2019; Zhou et al.,, 2022) only utilize the optical flow field between the onset frame and the apex frame. Therefore, SKD-TSTSAN can leverage motion information from different temporal regions for temporal modeling, fully considering the temporal information of ME and achieving the best recognition performance.
The performance of SKD-TSTSAN on the MMEW dataset is compared with different methods in Table 3. We conduct experiments on the MMEW dataset using CapsuleNet (Van Quang et al.,, 2019), MMNet (Li et al., 2022a, ), STSTNet (Liong et al.,, 2019), and TSCNN (Song et al.,, 2019). For MMNet’s performance on the 4-category MER task, the optimal performance reported in their original paper is provided. For six categories, CapsuleNet obtains the lowest UF1 and UAR since it only uses the apex frame as the network
@@
FR (Zhou et al.,, 2022) 3 34.93 34.13
STSTNet (Liong et al.,, 2019) 3 37.95 37.92
RCN-A (Xia et al.,, 2020) 3 39.28 38.93
$\mu$-BERT (Nguyen et al.,, 2023) 3 56.04 61.25
@@
AlexNet (Li et al., 2022b, ) 4 29.15 29.10
AlexNet(+Depth) (Li et al., 2022b, ) 4 30.01 29.82
SFAMNet (Liong et al.,, 2024) 4 44.62 47.97
$\mu$-BERT (Nguyen et al.,, 2023) 4 47.18 49.13
@@
AlexNet (Li et al., 2022b, ) 7 17.59 18.01
AlexNet(+Depth) (Li et al., 2022b, ) 7 17.73 18.29
SFAMNet (Liong et al.,, 2024) 7 23.65 23.73
$\mu$-BERT (Nguyen et al.,, 2023) 7 32.64 32.54
@@
input, which makes it difficult to extract the motion and temporal information of MEs. In contrast, STSTNet and TSCNN use the optical flow as input and achieve better performance. Our SKD-TSTSAN achieves the best recognition performance, reaching 77.11% UF1 and 74.77% UAR and showing improvements of 6.68% and 6.01% over TSCNN. Similar improvements are observed in the 4-category emotion classification: a 3.95% over TSCNN for UF1 and a 2.99% increase over MMNet for UAR.
To further validate our proposed SKD-TSTSAN, we perform comparative experiments on the large-scale ME dataset CAS(ME)3 using 3, 4, and 7 emotion categories, as shown in Table 4. When tested with 3, 4, and 7 emotion categories, SKD-TSTSAN exhibits double-digit improvements over all other methods. In the case of 3 emotion categories, SKD-TSTSAN obtains a UF1 score of 86.48% and a UAR score of 83.44%, which indicates an improvement of 30.44% in UF1 and 22.19% in UAR compared to the second-ranked $\mu$-BERT (Nguyen et al.,, 2023). For the 4 emotion categories, our SKD-TSTSAN maintains a considerable advantage over other methods with over 29.22% in UF1 and 23.44% in UAR. Using all 7 emotion categories is more demanding on the recognition performance of the MER method. SKD-TSTSAN significantly outperforms $\mu$-BERT in terms of UF1 (59.95% vs 32.64%) and UAR (57.31% vs 32.54%).
@@
datasets, SKD-TSTSAN achieves good recognition results for all categories. On the MMEW dataset, SKD-TSTSAN performs well on the 4-class MER task. On the 6-class MER task, SKD-TSTSAN has poor recognition performance for the “Fear” and “Sadness” samples. One of the main reasons is the small number of samples for the “Fear” and “Sadness” emotion categories. Besides, different emotion categories may correspond to the same AU. For example, both “Fear” and “Surprise” expressions appear in AU5 (“Upper Lid Raiser”), while “Fear” and “Disgust” expressions both involve AU4 (“Brow Lowerer”), as shown in Fig. 6. Therefore, samples labeled as “Fear” are easily misclassified as “Surprise” and “Disgust” categories, which have larger sample sizes. On the CAS(ME)3 dataset, the performance of SKD-TSTSAN is also affected by the imbalanced class distribution. For the 3 and 4 emotion categories, SKD-TSTSAN has poorer recognition performance on “Positive”, which has fewer samples. For the 7 emotion categories, SKD-TSTSAN performs better in the “Disgust”, “Others”, and “Surprise” categories. One of the reasons for the poor performance on the CAS(ME)3
@@
dataset is that the onset frame annotations of some samples are the same as the apex frame annotations. This results in no motion information in the optical flow between them, leading to misclassification.